40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed May 25 16:56:44 2022, Paco, Configuration, BHD, IFO recovery - IMC alignment 
    Reply  Wed May 25 17:34:47 2022, yuta, Configuration, BHD, IFO recovery - IFO alignment Screenshot_2022-05-25_17-47-57.png
       Reply  Thu May 26 19:55:43 2022, yuta, Configuration, BHD, Oplevs centered, BHD DCPDs are now online elog_1Y2.JPGelog_BHD.JPGelog_box.JPGScreenshot_2022-05-26_17-37-27_IFOaligned_OplevCentered.pngScreenshot_2022-05-26_20-35-02.png
          Reply  Fri May 27 17:45:53 2022, yuta, Configuration, BHD, BHD camera installed, GRY aligned 
             Reply  Thu Jun 2 20:05:37 2022, yuta, Configuration, PSL, IMC input power recovered to 1W, some alignment works Before.JPGAfter.JPG
                Reply  Fri Jun 3 12:13:58 2022, Paco, Configuration, CDS, Fix RFM channels  SoGreen.png
                   Reply  Fri Jun 3 15:22:51 2022, yuta, Update, LSC, Both arms locked with POY/POX, IR beam centered on TMs with ASS IRBeamsOnTMs.JPGScreenshot_2022-06-03_15-03-51.png
                      Reply  Fri Jun 3 17:42:50 2022, yuta, Update, LSC, MICH locks with AS55_Q Screenshot_2022-06-03_17-41-55.png
                         Reply  Sun Jun 5 19:46:40 2022, Paco, Update, LSC, Fixed IMC Trans sum issue 
                         Reply  Mon Jun 6 13:35:11 2022, Paco, Update, LSC, First calibrated spectra of MICH at AS55 Q Screenshot_2022-06-06_17-30-16_MICHOLTF.pngCalibrated_MICH_ERR.pdf
                            Reply  Fri Jun 17 16:22:21 2022, yuta, Update, LSC, Actuator calibration of BS. ITMX, ITMY, updated MICH displacement spectra from c1cal 6x
                               Reply  Thu Jul 7 18:18:19 2022, yuta, Update, LSC, Actuator calibration of ETMX and ETMX Screenshot_2022-07-05_14-52-01_OLTF.pngScreenshot_2022-07-05_14-54-03_TF.pngScreenshot_2022-07-05_14-56-41_Ratio.png
                                  Reply  Thu Jul 7 18:22:12 2022, yuta, Update, LSC, Actuator calibration of MC2 using Yarm TF.pngMC2.png
                                     Reply  Fri Jul 8 16:18:35 2022, rana, Update, LSC, Actuator calibration of MC2 using Yarm 
                                     Reply  Mon Jul 18 17:07:12 2022, yuta, Update, LSC, x4.12 added to ETMX coil outputs to balance with ETMY 
          Reply  Thu Jun 16 18:23:15 2022, Paco, Configuration, BHD, Recovering LO beam in BHD DCPDs Screenshot_2022-06-16_18-29-14_BHDLObeamISBACK.png
          Reply  Tue Jun 21 14:17:50 2022, yuta, Configuration, BHD, BHD DCPDs re-routed to c1sus2 C1X07ADC1.JPGBHDDCPDs.JPG
             Reply  Tue Jul 19 16:00:34 2022, yuta, Configuration, BHD, Fast channels for BHD DCPDs now available in c1lsc but not in c1hpc Screenshot_2022-07-19_14-26-39_c1hpc.pngScreenshot_2022-07-19_14-24-49_c1lsc.pngScreenshot_2022-07-19_15-51-25_GreenGreen.png
                Reply  Thu Jul 21 21:50:47 2022, Tega, Configuration, BHD, c1sus2 IPC update 
                   Reply  Fri Jul 22 15:05:26 2022, Tega, Configuration, BHD, c1sus2 shared memory and ADC fix Screenshot_2022-07-22_15-04-33_DCPD.pngScreenshot_2022-07-22_15-12-19_models.pngScreenshot_2022-07-22_15-15-11_ERR.pngScreenshot_2022-07-22_15-32-19_GDS.png
                      Reply  Fri Jul 22 17:46:10 2022, yuta, Configuration, BHD, c1sus2 watchdog update and DCPD ERR channels Screenshot_2022-07-22_17-48-25.png
                         Reply  Mon Jul 25 17:58:10 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update Screen_Shot_2022-07-25_at_5.43.28_PM.pngScreen_Shot_2022-07-25_at_5.21.10_PM.pngScreen_Shot_2022-07-25_at_5.30.58_PM.pngScreen_Shot_2022-07-25_at_5.35.19_PM.png
                            Reply  Mon Aug 1 18:42:39 2022, Tega, Configuration, BHD, c1sus2 IPC dolphin issue update c1sus2_dolphin.pngfb1_dxamin_status.pngdolphin_num_mem_init2.png
                               Reply  Tue Aug 2 17:25:18 2022, Tega, Configuration, BHD, c1sus2 dolphin IPC issue solved c1lsc_IPC_status.pngFE_lsmod_dependencies_c1sus2_b4_after_iop_unpdate.png
                      Reply  Mon Jul 25 18:09:41 2022, Tega, Configuration, BHD, BHD Homodyne Phase control MEDM screen Screen_Shot_2022-07-25_at_6.12.08_PM.pngScreen_Shot_2022-07-25_at_6.18.09_PM.png
Message ID: 17018     Entry time: Tue Jul 19 16:00:34 2022     In reply to: 16932     Reply to this: 17025
Author: yuta 
Type: Configuration 
Category: BHD 
Subject: Fast channels for BHD DCPDs now available in c1lsc but not in c1hpc 

[Paco, Anchal-remote-support, Yuta]

We added fast channels to BHD DC PDs.
C1:LSC-DCPD_(A|B)_IN1 are now available, but C1:HPC-DCPD_(A|B)_IN1 still gives us zero.

c1hpc situation -> not good
 - We can see the slow signal at C1:X07-MADC1_EPICS_CH16 (DC PD A) and CH17 (DC PD B)
 - C1:HPC-DCPD_(A|B)_IN1 is there, but zero.
 - We have modified c1hpc model to add DCPD_(A|B) filters in front of the input matrix (see Attachment #1).
 - After modifying the model, we run
ssh c1sus2
rtcds make c1hpc
rtcds install c1hpc
ssh fb1
sudo systemctl restart daqd_*

 - After this, we got 0x2000 error. So, we ran the following. This removed 0x2000 error, but DCPD signals are still zero. They are also not available in C1HPC-MONITOR_ADC1.adl screen (see Attachment #3).
ssh c1sus2
rtcds restart c1hpc


c1lsc situation -> good
 - We could see the slow signal at C1:X04-MADC1_EPICS_CH4 (DC PD A) and CH5 (DC PD B), and also C1:LSC-DCPD_(A|B)_NORM after making C1:LSC-DCPD_(A|B)_POW_NORM=1. The ADC channel and DCPD channel are exactly the same.
 - After confirming the above, we modified the c1lsc model to add DCPD_(A|B) filters in front of the input matrix (see Attachment #2).
 - After modifying the model, we run
ssh c1lsc
rtcds make c1lsc
rtcds install c1lsc
ssh fb1
sudo systemctl restart daqd_*

 - After this, we also got 0x2000 error. We also noticed that, for example, C1:X04-MADC0_EPICS_CH31 and C1:LSC-ASDC_INMON are different, which used to be the same (ASDC_INMON was largely attenuated).
 - In the end, we run the following to remove 0x2000 error, but it crashed c1lsc, as well as c1sus, c1ioo.
ssh c1lsc
rtcds restart c1lsc

 - So, we did rebootC1LSC.sh. This made c1lsc, c1ioo and c1sus as green as before, except for RFM issue in TRX/TRY, like we saw in June. We followed the steps in 40m/16887 to hard reboot c1iscex/c1iscey and ran rebootC1LSC.sh again. This made C1CDS_FE_STATUS.adl screen as green as before (see Attachment #3).

 - Fast channels C1:LSC-DCPD_(A|B)_IN1 are now available. They are also available in C1LSC-MONITOR_ADC1.adl screen (see Attachment #3).

Attachment 1: Screenshot_2022-07-19_14-26-39_c1hpc.png  61 kB  Uploaded Tue Jul 19 17:02:02 2022  | Hide | Hide all
Screenshot_2022-07-19_14-26-39_c1hpc.png
Attachment 2: Screenshot_2022-07-19_14-24-49_c1lsc.png  152 kB  Uploaded Tue Jul 19 17:03:30 2022  | Hide | Hide all
Screenshot_2022-07-19_14-24-49_c1lsc.png
Attachment 3: Screenshot_2022-07-19_15-51-25_GreenGreen.png  57 kB  Uploaded Tue Jul 19 17:05:36 2022  | Hide | Hide all
Screenshot_2022-07-19_15-51-25_GreenGreen.png
ELOG V3.1.3-